### v1

@@ III Conclusion and discussions
First, it will be insightful to understand our results in the context of the decorated domain wall construction. For fermionic SPT phases, the construction produces several “layers” of decorations Wang and Gu (2020). In (4+1)d, for a symmetry group $G$, the decorations are $p_{x}+\hskip 1.0pt\mathrm{i}\hskip 1.0ptp_{y}$ superconductors on 2D defects (classified by $\operatorname{H}^{2}(G,\mathbb{Z})$), Majorana chains on 1D defects (classified by $\operatorname{H}^{3}(G,\mathbb{Z}_{2})$), and complex fermions on 0D defects (classified by $\operatorname{H}^{4}(G,\mathbb{Z}_{2})$). One can show that the $\nu=N$ state is characterized by the Majorana chain decoration, while those with $p_{x}+\hskip 1.0pt\mathrm{i}\hskip 1.0ptp_{y}$ decorations do not admit a gapped boundary TQFT description. A natural conjecture is that the same is true for other symmetry groups in (3+1)d as well. Proving or disproving this conjecture is an interesting question for future work.
In general, topological boundary states for FSPT phases with Majorana chain decorations are not well-understood. In (2+1)d, the only known case is the semion-fermion theory, which carries $\nu=2$ anomaly (Majorana chain decoration) of the (3+1)d topological superconductor. However, for Majorana FSPTs in (3+1)d protected by unitary finite-group symmetry, no explicit examples of topological boundary states are known. These questions are special cases under a broader question of characterizing ’t Hooft anomalies in topological orders enriched by global symmetries. Namely, given a TQFT and how symmetry acts in the theory, how can we compute the ’t Hooft anomaly associated with the symmetry using the given data. While this problem has been extensively studied in (2+1)d (see e.g. Barkeshli and Cheng (2020); Tata et al. (2022); Bulmash and Barkeshli (2022)), the (3+1)d case has been much less explored, which may be a fruitful avenue for future research.
@@
Acknowledgement
MC is supported by NSF grant DMR-2424315. The authors are listed alphabetically by a standard convention.
@@ Acknowledgement
$\begin{cases}\mathbb{Z}_{N}&\quad N\equiv 1\,(\text{mod }2)\\ \mathbb{Z}_{4N}=\mathbb{Z}_{8}\oplus\mathbb{Z}_{N/2}&\quad N\equiv 2\,(\text{mod }4)\\ \mathbb{Z}_{2N}\oplus\mathbb{Z}_{2}&\quad N\equiv 0\,(\text{mod }4)\end{cases}\,.$ (16)
@@
$(\nu,m,\lambda),\,\nu\in\mathbb{Z},\,m\in\mathbb{Z}_{2N},\,\lambda\in\mathbb{Z}_{2}\,,$ (17)
where the index $\nu$ labels the $\mathbb{Z}$ classification of $p+ip$, $\nu$ labels the order of the topological spin of the $\mathbb{Z}_{N}$ flux and $\lambda$ indicates the existence of MZM. We start with a topologically trivial block state, consisting of $N$ layers of $p+ip$ superconductors, on which the $C_{N}$ acts as the $\mathbb{Z}_{N}$ cyclic permutation of the layers.
@@
The trivial block state corresponds to $(N,m_{N},1)$, thus we find that $(-N,0,0)\sim(0,m_{N},1)$ 777$\sim$ means phase equivalence.. Denote $r_{N}=\frac{2^{2p-1}+1}{3}$, which is easily seen to be an odd integer. This indicates that $(0,m_{N},0)$ has order $8$ in $\mathbb{Z}_{2N}=\mathbb{Z}_{2^{p+1}}$. Therefore, the element $(1,0,0)$ has order $8N=2^{p+3}$.
@@
    $\begin{cases}\mathbb{Z}_{N}\oplus\mathbb{Z}_{N}&3\nmid N\\ \mathbb{Z}_{3N}\oplus\mathbb{Z}_{N/3}&3\;|\;N\end{cases}\,;$ (21)
@@ Appendix B $\mathbb{Z}_{16}$ class anomaly of $\mathbb{Z}_{4}^{\rm F}$ symmetry in the Standard Model
    $15\times 3$ Weyl fermions and an anomaly index $-3+n_{\nu_{R}}$:
    The SM has $15$ Weyl fermions per family (or per generation), and with $N_{f}=3$ families thus a total $15\times 3=45$ Weyl fermions of quarks and leptons. We do not yet know how many types of right-handed neutrinos there are beyond the SM: the total number of types of right-handed neutrinos, say $n_{\nu_{R}}$, can be $n_{\nu_{R}}=0,1,2,3,\dots$ and so on. The potentially nonzero number $-N_{f}+n_{\nu_{R}}=-3+n_{\nu_{R}}$ is an important anomaly index for the SM.
@@
    This specific combination as an order four finite cyclic group $\mathbb{Z}_{4,X\equiv 5({\bf B}-{\bf L})-\frac{2}{3}\tilde{Y}}$ is also pointed out by Wilczek-Zee Krauss and Wilczek (1989b). The $\mathbb{Z}_{4,X}$ symmetry in the SM is exactly the $\mathbb{Z}_{4}^{\rm F}$ symmetry discussed in our main text. Recently, in [Garcia-Etxebarria and Montero, 2019; Wan and Wang, 2020; Wang, 2020a, b, 2021; Wang et al., 2022; Putrov and Wang, 2023], it was pointed out that the (3+1)d SM suffers from a mod 16 class of this mixed $\mathbb{Z}_{4,X}$-gauge-gravity nonperturbative global anomaly captured by a (4+1)d fermionic invertible topological field theory (namely, the low-energy field theory of the (4+1)d FSPT that we studied)
@@
    $\psi_{q}\psi_{q}\psi_{q}\psi_{l},\quad\psi_{\bar{q}}\psi_{q}\psi_{\bar{l}}\psi_{l},\quad\psi_{\bar{q}}\psi_{q}\psi_{\bar{q}}\psi_{q},\quad\psi_{\bar{l}}\psi_{l}\psi_{\bar{l}}\psi_{l}.$ (26)
@@
    As mentioned in the introduction, the $\mathbb{Z}_{16}$ class of the $\mathbb{Z}_{4}^{\rm F}$ chiral anomaly shows up as the mixed gauge-gravity nonperturbative global anomaly for $\mathbb{Z}_{4,X}$ in the SM. Three generations of experimentally confirmed quarks and leptons contribute a total number of Weyl fermions in the SM $3\times 15=45$, which gives an anomaly index $45\mod 16=-3\mod 16$. The $-3\mod 16$ is the anomaly index of the SM, for the $\mathbb{Z}_{16}$ class of the $\mathbb{Z}_{4,X}$-gauge-gravitational nonperturbative global anomaly. As a speculation, the (3+1)d boundary topological order that we construct together with the (4+1)d bulk FSPT provides an alternative way to cancel the $-3\mod 16$ class of anomaly in the SM, without necessarily including 3 generations of right-handed neutrinos. The experimental absence of any right-handed neutrinos in the SM provides us the theoretical opportunity to include the exotic topological BSM sector via appending the (3+1)d non-invertible topological order or (4+1)d invertible topological phase to the SM. For example, anomaly cancellation can be achieved by having one right-handed neutrino, together with the (3+1)d $\mathbb{Z}_{4}$ gauge theory TQFT discussed here. Thus, these topological states (and the fractionalized/anyonic energetic excitations above the topological order energy gap) become a candidate for the BSM cold dark matter Wang (2020a, b, 2021).

### v2

@@ III Conclusion and discussions
First, it will be insightful to understand our results in the context of the decorated domain wall construction. For fermionic SPT phases, the construction produces several “layers” of decorations Wang and Gu (2020). In (4+1)d, for a symmetry group $G$, the decorations are $p_{x}+\hskip 1.0pt\mathrm{i}\hskip 1.0ptp_{y}$ superconductors on 2D defects (classified by $\operatorname{H}^{2}(G,\mathbb{Z})$), Majorana chains on 1D defects (classified by $\operatorname{H}^{3}(G,\mathbb{Z}_{2})$), and complex fermions on 0D defects (classified by $\operatorname{H}^{4}(G,\mathbb{Z}_{2})$). One can show that the $\nu=N$ state is characterized by the Majorana chain decoration, while those with $p_{x}+\hskip 1.0pt\mathrm{i}\hskip 1.0ptp_{y}$ decorations do not admit a gapped boundary TQFT description. A natural conjecture is that the same is true for the boundary topological orders of (4+1)d fermionic SPT phase with other symmetry groups. Proving or disproving this conjecture is an interesting question for future work.
We now give an algebraic argument that the $p+\hskip 1.0pt\mathrm{i}\hskip 1.0ptp$ layer cannot be trivialized by symmetry extension, i.e. the 5-dimensional term
$A\cup p_{1}\equiv p_{1}({\rm PD}(A)),$
where $A\in\operatorname{H}^{1}(G,{\rm U}(1))$ paired with the Pontryagin class $p_{1}$, cannot be trivialized by symmetry extension for $G$ being a finite group. Given that all (3+1)d TQFTs are gauged finite group SPTs, this argument strongly suggests that no symmetric boundary TQFT exists for (4+1)d fermionic SPTs with nontrivial $p+\hskip 1.0pt\mathrm{i}\hskip 1.0ptp$ decoration.
The $p+\hskip 1.0pt\mathrm{i}\hskip 1.0ptp$ decoration is classified by $\operatorname{H}^{2}(G,\mathbb{Z})$. Since $G$ is finite, $\operatorname{H}^{2}(G,\mathbb{Z})$ is isomorphic to $\operatorname{H}^{1}(G,{\rm U}(1))$, i.e. group of one-dimensional representations of $G$. Given a nontrivial $\omega\in\operatorname{H}^{2}(G,\mathbb{Z})$, we have a nontrivial 1D representation $\lambda\in\operatorname{H}^{1}(G,{\rm U}(1))\cong\operatorname{Hom}(G,{\rm U}(1))$ ($\omega$ is related to $\lambda$ via the Bockstein homomorphism). Consider a group extension of $G$ by a finite group $K$:
$1\rightarrow K\rightarrow G^{\prime}\stackrel{{\scriptstyle\mathrm{r}}}{{\rightarrow}}G\rightarrow 1,$
where the surjective map from $G^{\prime}$ to $G$ is denoted by a reduction $\mathrm{r}$. Suppose $\lambda$ is lifted to $\lambda^{\prime}\in\operatorname{H}^{1}(G^{\prime},{\rm U}(1))$, then the pullback requires $\lambda^{\prime}(g^{\prime})=(\mathrm{r}^{*}\lambda)(g^{\prime}):=\lambda(\mathrm{r}(g^{\prime}))$ for $g^{\prime}\in G^{\prime}$. As $\mathrm{r}^{*}$ is injective on $\operatorname{H}^{1}(-,{\rm U}(1))$, a nontrivial $\lambda$ lifts to a nontrivial $\lambda^{\prime}$, giving a nontrivial class in $\operatorname{H}^{2}(G^{\prime},\mathbb{Z})$. Therefore, $\omega\in\operatorname{H}^{2}(G,\mathbb{Z})$ cannot be trivialized by symmetry extension.
In general, topologically ordered boundary states for FSPT phases with Majorana chain decorations are not well-understood. In (2+1)d, the only known case is the semion-fermion theory, which carries $\nu=2$ anomaly (Majorana chain decoration) of the (3+1)d topological superconductor with time-reversal symmetry. However, for Majorana FSPTs in (3+1)d protected by unitary finite-group symmetry, no explicit examples of topological boundary states are known. These questions are special cases of a broader problem of characterizing ’t Hooft anomalies of topological orders enriched by global symmetries. Namely, given a TQFT and how symmetry acts in the theory, how can we compute the ’t Hooft anomaly associated with the symmetry using the given data? While this problem has been extensively studied in (2+1)d (see e.g. Barkeshli and Cheng (2020); Tata et al. (2022); Bulmash and Barkeshli (2022)), the (3+1)d case has been much less explored, which may be a fruitful avenue for future research.
@@
Acknowledgment
MC is supported by NSF grant DMR-2424315. JW is supported by LIMS and Ben Delo Fellowship. The authors are listed alphabetically.
Note added: After the completion of this work, [Debray et al., 2025] and [Wan and Wang, 2025] appeared, reproducing the result of this work using the symmetry extension approach.
@@ Acknowledgment
$\begin{cases}\mathbb{Z}_{N}&\quad N\equiv 1\,(\text{mod }2)\\ \mathbb{Z}_{4N}=\mathbb{Z}_{8}\oplus\mathbb{Z}_{N/2}&\quad N\equiv 2\,(\text{mod }4)\\ \mathbb{Z}_{2N}\oplus\mathbb{Z}_{2}&\quad N\equiv 0\,(\text{mod }4)\end{cases}\,.$
@@
$(\nu,m,\lambda),\,\nu\in\mathbb{Z},\,m\in\mathbb{Z}_{2N},\,\lambda\in\mathbb{Z}_{2}\,$,
where the index $\nu$ labels the $\mathbb{Z}$ classification of $p+ip$, $m$ labels the order of the topological spin of the $\mathbb{Z}_{N}$ flux and $\lambda$ indicates the existence of MZM. We start with a topologically trivial block state, consisting of $N$ layers of $p+ip$ superconductors, on which the $C_{N}$ acts as the $\mathbb{Z}_{N}$ cyclic permutation of the layers.
@@
$e^{\frac{\hskip 1.0pt\mathrm{i}\hskip 1.0pt\pi m}{N}}={\Theta}=e^{\frac{\hskip 1.0pt\mathrm{i}\hskip 1.0pt\pi}{24}(N^{2}+2)}\,\;\Longrightarrow\;\,m_{N}=\frac{N(N^{2}+2)}{24}=2^{p-2}\frac{2^{2p-1}+1}{3}.$ (17)
@@
    $\begin{cases}\mathbb{Z}_{N}\oplus\mathbb{Z}_{N}&3\nmid N\\ \mathbb{Z}_{3N}\oplus\mathbb{Z}_{N/3}&3\;|\;N\end{cases}\,;$
@@ Appendix B $\mathbb{Z}_{16}$ class anomaly of $\mathbb{Z}_{4}^{\rm F}$ symmetry in the Standard Model
    $15\times 3$ Weyl fermions and an anomaly index $3+n_{\nu_{R}}$:
    The SM has $15$ Weyl fermions per family (or per generation), and with $N_{f}=3$ families thus a total $15\times 3=45$ Weyl fermions of quarks and leptons. We do not yet know how many types of right-handed neutrinos there are beyond the SM: the total number of types of right-handed neutrinos, say $n_{\nu_{R}}$, can be $n_{\nu_{R}}=0,1,2,3,\dots$ and so on. The potentially nonzero number $N_{f}+n_{\nu_{R}}=-3+n_{\nu_{R}}$ is an important anomaly index for the SM.
@@
    This specific combination as an order four finite cyclic group $\mathbb{Z}_{4,X\equiv 5({\bf B}-{\bf L})-\frac{2}{3}\tilde{Y}}$ is also pointed out by Wilczek-Zee Krauss and Wilczek (1989b). The $\mathbb{Z}_{4,X}$ symmetry in the SM is exactly the $\mathbb{Z}_{4}^{\rm F}$ symmetry discussed in our main text. Recently, in [García-Etxebarria and Montero, 2019; Wan and Wang, 2020; Wang, 2020a, b, 2021; Wang et al., 2022a, b; Putrov and Wang, 2024], it was pointed out that the (3+1)d SM suffers from a mod 16 class of this mixed $\mathbb{Z}_{4,X}$-gauge-gravity nonperturbative global anomaly captured by a (4+1)d fermionic invertible topological field theory (namely, the low-energy field theory of the (4+1)d FSPT that we studied)
@@
    $\psi_{q}\psi_{q}\psi_{q}\psi_{l},\quad\psi_{\bar{q}}\psi_{q}\psi_{\bar{l}}\psi_{l},\quad\psi_{\bar{q}}\psi_{q}\psi_{\bar{q}}\psi_{q},\quad\psi_{\bar{l}}\psi_{l}\psi_{\bar{l}}\psi_{l}.$
@@
    As mentioned in the introduction, the $\mathbb{Z}_{16}$ class of the $\mathbb{Z}_{4}^{\rm F}$ chiral anomaly shows up as the mixed gauge-gravity nonperturbative global anomaly for $\mathbb{Z}_{4,X}$ in the SM. Three generations of experimentally confirmed quarks and leptons contribute a total number of Weyl fermions in the SM $3\times 15=45$, which gives an anomaly index $45\mod 16=-3\mod 16$. The $3\mod 16$ is the anomaly index of the SM, for the $\mathbb{Z}_{16}$ class of the $\mathbb{Z}_{4,X}$-gauge-gravitational nonperturbative global anomaly. As a speculation, the (3+1)d boundary topological order that we construct together with the (4+1)d bulk FSPT provides an alternative way to cancel the $3\mod 16$ class of anomaly in the SM, without necessarily including 3 generations of right-handed neutrinos. The experimental absence of any right-handed neutrinos in the SM provides us the theoretical opportunity to include the exotic topological BSM sector via appending the (3+1)d non-invertible topological order or (4+1)d invertible topological phase to the SM. For example, anomaly cancellation can be achieved by having one right-handed neutrino, together with the (3+1)d $\mathbb{Z}_{4}$ gauge theory TQFT discussed here. Thus, these topological states (and the fractionalized/anyonic energetic excitations above the topological order energy gap) become a candidate for the BSM cold dark matter Wang (2020a, b, 2021).